I am trying to find the absolute values of the equation -2[3j] - 8 <= -20. i have ran through it numerous times and cant seem to get it right.
Answer by tommyt3rd(5050)

-2[3j] -8 <= -20
-2[3j] -8+8 <= -20+8
-2[3j] <= -12
-2[3j]/(-2) >= 6

[3j]>=6

[j] >=2


so

-j<= -2 or j>=2

:)